SpaceRef

SpaceRef


NASA Hubble Space Telescope Daily Status Report #4715

Status Report From: Space Telescope Science Institute
Posted: Wednesday, October 15, 2008

HUBBLE SPACE TELESCOPE DAILY REPORT #4715

Continuing to collect World Class Science

PERIOD COVERED: 5am October 10 - 5am October 14, 2008 (DOY 284/0900z-288/0900z)

OBSERVATIONS SCHEDULED

FGS 11212

Filling the Period Gap for Massive Binaries

The current census of binaries among the massive O-type stars is seriously incomplete for systems in the period range from years to millennia because the radial velocity variations are too small and the angular separations too close for easy detection. Here we propose to discover binaries in this observational gap through a Faint Guidance Sensor SNAP survey of relatively bright targets listed in the Galactic O Star Catalog. Our primary goal is to determine the binary frequency among those in the cluster/association, field, and runaway groups. The results will help us assess the role of binaries in massive star formation and in the processes that lead to the ejection of massive stars from their natal clusters. The program will also lead to the identification of new, close binaries that will be targets of long term spectroscopic and high angular resolution observations to determine their masses and distances. The results will also be important for the interpretation of the spectra of suspected and newly identified binary and multiple systems.

FLIGHT OPERATIONS SUMMARY:

Significant Spacecraft Anomalies: (The following are preliminary reports of potential non-nominal performance that will be investigated.)

HSTARS:

11529 - GSAcq (2,0,2) Failed to RGA Hold (Gyro Control) GSAcq(2,0,2) scheduled at 287/00:48:27 - 287/00:54:44 failed to RGA Hold. Acquisition summary revealed FGS-2 achieved fine lock then transitioned to SSM control at 287/00:54:03.

Observation affected: Astrometry only, Proposal #11212.

11530 - GSAcq(2,0,2) failed to RGA Hold (Gyro Control) Upon acquisition of signal at 288/05:06:40, the GSAcq(2,0,2)scheduled at 288/03:52:34 - 03:58:51 had failed to RGA Hold due to (QF2STOPF) stop indication flag set on FGS-2.

Observation affected: Astrometry only, Proposal ID 11212.

COMPLETED OPS REQUEST:

18307-0 - Restore HV Protection for the SSM @ 284/14:25z

COMPLETED OPS NOTES: (None)

                        SCHEDULED      SUCCESSFUL 
FGS GSacq               39                  37 
FGS REacq               00                  00 
OBAD with Maneuver      78                  78 

SIGNIFICANT EVENTS:

FLASH: Restoration of HV Protection

At 284/14:25:46 UTC, the SSM HV Protection capability through the NSSC-1 was successfully restored via Ops Request 18307-0. Executive RTCS 21 was enabled to send notification of an HV Protect event through the setting of the former Mechanism Motion flag, and NSSC-1 Global Event Flag 12 was set to allow this transfer across the PIT to the SSM.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

NanoRacks